OpenStudy (anonymous):

Expand and simplify (√3+2)^5. Give your answer in the form a+b√3

OpenStudy (dumbcow):

use binomial thm: \[(a+b)^{5} = (5C0) a^{5} + (5C1) a^{4}b+(5C2) a^{3}b^{2}+(5C3)a^{2}b^{3}+(5C4)a b^{4}+(5C5)b^{5}\]

OpenStudy (anonymous):

what is \[(5C)?\]

OpenStudy (dumbcow):

sorry that is notation for combinations 5 choose 0 = 5!/0!5! 5 choose 3 = 5!/3!2!

OpenStudy (dumbcow):

i assumed you have learned about combinations...you can use a calculator to compute them as well

OpenStudy (anonymous):

5!/0!5! = 1

OpenStudy (dumbcow):

yes 5C0 = 1 5C1 = 5 5C2 = 10 5C3 = 10 5C4 = 5 5C5 = 1

OpenStudy (anonymous):

\[(\sqrt{3})^{5}+5(\sqrt{3})^{4}2+10(\sqrt{3})^{3}2^{2}+10(\sqrt{3})^{2}2^{3}+5(\sqrt{3})2^{4}+2^{5}\]

OpenStudy (dumbcow):

yep, now hard part....simplify sqrt3 * sqrt3 = 3

OpenStudy (anonymous):

oh god ok

OpenStudy (dumbcow):

to check your answer...i gotta go http://www.wolframalpha.com/input/?i=%28sqrt%283%29%2B2%29%5E5